OpenStudy (anonymous):

What is the value of x in the solution to the following system of equations? x − y = −3 x + 3y = 5

OpenStudy (anonymous):

@Hero @tHe_FiZiCx99

myininaya (myininaya):

elimination style: to set the y's up for elimination multiply first equation by 3 then add the equations together and tell me what you get

OpenStudy (anonymous):

I dont know how to do that?

myininaya (myininaya):

multiplying or adding?

OpenStudy (anonymous):

2x - 2y= 2 ?

myininaya (myininaya):

The first equation is x-y=-3 I asked you to multiply the first equation by 3. 3(x-y=-3)

myininaya (myininaya):

3(x-y)=3(-3) can you distribute the 3 on the left hand side and multiply 3 and -3 on the right hand side

OpenStudy (anonymous):

So what's the answer?

myininaya (myininaya):

can you do what I asked and you will get your answer hang in there

OpenStudy (anonymous):

3(x-y)? and -9?